OpenStudy (anonymous):

The mapping x to x^2 from a group G to itself is an automorphism. What can you say about G? Assume G is finite.

OpenStudy (anonymous):

I found out G is abelian, using the fact that the map is a homomorphism. And that no non-identity element has order 2. I cant proceed further. :(

terenzreignz (terenzreignz):

hmm... let's see :)

terenzreignz (terenzreignz):

Automorphisms are isomorphisms?

OpenStudy (anonymous):

Yes

terenzreignz (terenzreignz):

Maybe cyclic?

terenzreignz (terenzreignz):

Probably not, though, there's no element that may even remotely suggest a generator...

OpenStudy (anonymous):

I found out all cyclic groups of odd order have this automorphism through trial and error. How do I prove that G is such a group only?

terenzreignz (terenzreignz):

We don't even know if its order is finite yet... :D

OpenStudy (anonymous):

Oops. The question also assumes that G is finite. Its mentioned! My bad. Proceed using finite order group.

terenzreignz (terenzreignz):

Okay, so maybe it's cyclic... let's find out :D

OpenStudy (anonymous):

If its cyclic of even order, then it would have an non-identity element of order 2. So it cant be cyclic of even order. (since, injective so kernel is trivial)

OpenStudy (anonymous):

How do we prove G is cyclic?

terenzreignz (terenzreignz):

I don't even know if it is :) I'm just shooting random ideas about what we could possibly conclude about G :)

terenzreignz (terenzreignz):

Any other ideas about what we can conclude about the group? There are so few special groups :D

OpenStudy (anonymous):

What I am thinking is that, x is mapped to x^2 which is mapped to x^4 which is mapped to x^8 and so on. But this process must stop at some x^(2m) since G is finite.

terenzreignz (terenzreignz):

What does that tell us? :)

OpenStudy (anonymous):

That at some final stage, x^(2m)=x. That is, x^(2m-1)=1. So x has order (2m-1). So, every element in G has odd order. But I dont see how to correlate cyclicity and this!

terenzreignz (terenzreignz):

We don't know that x^2m = x do we? We just know that at some point, x^2m = x^2n where n < m

OpenStudy (anonymous):

Why? x is mapped to x^2 and x^2 to x^4 .... goes on... finally x^m mapped to x^2m, which cant be mapped further, cause there are no more powers of x as G is finite. So x^2m must be x. Am i correct?

terenzreignz (terenzreignz):

No... it just means we already passed x^2m at some point, but I don't see how it guarantees that it'd be x.

OpenStudy (anonymous):

Suppose all elements (powers of x) were never repeated until x^2m. Now x^2m cant be any other power of x than x. If it were, say, x^2m=x^2, then x^(2m-1) would have been the first element which was repeated, against our assumption.

terenzreignz (terenzreignz):

LOL okay, I'm convinced :)

terenzreignz (terenzreignz):

But still, that says nothing about whether or not it's cyclic... This stuff is hard without knowing what to aim for :D

OpenStudy (anonymous):

What we need to know now, is that are there any other elements other than powers of x in G? Is there a way to show that there isn't? Then we could prove that G is cyclic, hence killing the problem!

terenzreignz (terenzreignz):

Okay, let's start with that... suppose there is an element y such that \[\Large x^n \ne y\] for any positive integer n

terenzreignz (terenzreignz):

Wait, no... it might be cyclic, but how do we know that x is the generator?

OpenStudy (anonymous):

Yeah, but the map generates a subgroup of G, that is (x , x^2 , x^4 , ... , 1). Its a odd order subgroup and all it's elements are even powers. Can I tell that its <x>?

OpenStudy (anonymous):

Suppose order of x is 5. Then (x, x^2, x^4, x^8=x^3, 1). See, its a cyclic group.

OpenStudy (anonymous):

Is the answer all abelian groups of odd order OR is it all cyclic groups of odd order? What is an example of abelian group of odd order which is not cyclic?

terenzreignz (terenzreignz):

map generates a subgroup?